• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 08:05:02 on April 1,2005

    Entry number 142952

    keyword=Owl Shift Summary

    Owl Shift Summary
    ------------------
    
    01:35   MCC, per our request, increased current from 40 to 45 uA.
    
    02:04   Ended Compton Run (10527) and started a new one (10528).
    
    05:30   TS1 not reporting status, we had to reboot all the trigger supervisors 
    and restart CODA.  Run 3118 end of run failed.  Run 3119 is junk - just testing 
    the DAQ.  Also, the compton stripcharts had flatlined.  We had to reboot 
    compton.jlab.org and started a new compton run (10529).
    
    05:45   MCC called to say that vacuum was getting close to the trip point, so 
    we lowered our beam current to 40 uA.
    
    06:30   Noticed that compton rates were zero, so I ended run 10529 and started 
    a new run (10530).  Ok for now.
    
    07:00   Ended run 3120 because deadtime was high - 96%.  Rebooted TS2 and 
    everything seemed ok, but deadtime was still around 10%.  It turned out that 
    ROC5, id0 is not initialized. Tried reboot of TS2 again, but it didn't help.  
    Need to go to hall to reset VME crate.  Runs 3121-3124 -> junk.
    
    07:27	Called for controlled access to reboot VME crate.
    
    Bodo went to the hall for the reboot, and everything was successful.
    
    07:58:  MCC ready to restore beam to the hall.
    
    Production Runs on tilted Carbon
    
    Run   Start    End    # events
    _________________________________
    
    3114  00:00   00:54    233k
    3115  00:54   01:58    300k
    3116  02:01   03:06    300k
    3117  03:07   04:11    300k
    3118  04:11   05:00    220k
    3120  05:38   07:00    290k
    3125  07:59